DETERMINE WHETHER THE FOLLOWING STATEMENTS HAVE A KNOWN OR UNKNOWN POPULATION VARIANCE. IDENTIFY ALSO THE FORMULA TO BE USED TO ESTIMATE THE STANDARD ERROR OF THE MEAN.


An SHS teacher claims that the average time it takes a group of students to complete the Statistics and Probability examination is 50.5 minutes with a variance of 17.64 minutes squared. She randomly selected 45 students and found them to have a mean of 52 minutes and a standard deviation of 3.5 minutes. She then used the z-distribution to find out if the group can complete the exam faster than the population.


1
Expert's answer
2021-12-07T04:45:41-0500

We are given that,

"\mu=50.5,\space \sigma=17.64"

"n=45,\space \bar{x}=52,\space s=3.5"

Here, the population variance is known since we are provided with the population standard deviation.

The population variance is, "\sigma^2=(\sigma)^2=17.64^2=311.1696".

Since the population variance is known and the sample size is greater than 30, we apply the formula below to determine the standard error for the mean.

"S E=Z_{\alpha }*(\sigma/\sqrt{n})" where "Z_{\alpha}" is the standard normal table value at "\alpha" level of significance, "n" is the sample size and "\sigma" is the population standard deviation.

Therefore, the formula to estimate the standard error of the mean is "Z_{\alpha}(\sigma/\sqrt{n})"